A rectangular prism has a height of 4 millimeters and width 5 millimeterss. If the total surface area is 166 square millimeters, what is the length of the prism?
Answer by graphmatics(170)

if W = Width of the prism, L = Length of the prism and
H = Height of the prism, the surface area A of a rectangular prism is given by
A = 2WL + 2LH + 2HW
_____________________________
So as A = 166, H = 4 and W = 5
166 = 2*5*L + 2*4*L + 2*5*4
166 = 10*L + 8*L + 40
-18*L = -166 + 40
-18*L = -126
L = 7
So Length of the prism is 7 millimeters